# Scalable multi-qubit intrinsic gates in quantum dot arrays

### v1

@@ Abstract
We study the multi-qubit quantum gates intrinsic to a general array of semiconductor quantum dots and investigate how they can be implemented in a scalable way. The intrinsic quantum gates refer to the class of natural-forming transformations in the qubit rotating-frame under direct exchange coupling, and can be recognized as the instruction set of a spin-qubit chip. Adopting an perturbative treatment, we can model intrinsic gates by first-order dynamics in the coupling strength. A general formalism is developed for identifying the multi-qubit intrinsic gates under arbitrary array connectivity. Factors influencing the fidelities of the multi-qubit intrinsic gates are discussed. The advantageous applications of intrinsic gates in quantum computing and quantum error correction are explored. We also propose a theoretical scheme to overcome the problem of inhomogeneous coupling using dynamical calibration of the connecting bonds. This scheme can be further combined with periodic dynamical decoupling for robust implementations of multi-qubit gates in large-scale quantum computers.
@@ I Introduction
Semiconductor quantum dots are promising physical platforms for universal quantum computing [1, 2, 3, 4]. The spins of electrons (or holes) are natural two-level systems that can be selectively manipulated and brought into interactions by confining them with artificial structures. Spin qubits defined in quantum dots enjoy several unique advantages, including their small physical size and compatibility with modern semiconductor fabrication techniques, that make them suitable for creating large-scale quantum chips [5, 6]. Under intensive research efforts from the global community, the technologies and theories behind spin qubits have seen significant advancements over recent years. Continuous developments in key performance metrics such as the coherence times, operation frequencies and gate fidelities have been made [7]. Spin qubits in the state-of-the-art can be prepared and measured with fidelity exceeding $99\%$ [8], manipulated at frequency over 540 MHz [9] and work in temperature above 4K [10]. The fidelity figures for single-qubit and two-qubit gates have been pushed beyond the fault-tolerance threshold value [11, 12]. Industrial manufacturing have also been demonstrated for quantum dot arrays [13, 14].
@@
The natural roadmap of any quantum computing scheme is to incorporate gradually increasing quantities of qubits for realizing useful quantum algorithms. However, accurate control of multiple qubits can be a highly nontrivial task that implies much more than just putting the qubits together. For selective control of the interqubit coupling, one must be able to eliminate the crosstalk effects of the control signals in an efficient manner. In addition, there must be a robust method to overcome the inhomogeneous coupling across the qubits. The key challenge is that the relevant control resources must scale well with the number of qubits, for potential application to large-scale chips with millions of qubits. For spin qubits, a viable way to scale up the qubit count is to employ an increasing two-dimensional array of quantum dots, as recently demonstrated for crossbar arrays of hole qubits [15, 16, 17, 18]. This configuration is also compatible with the surface code, a topological error correction code commonly conceived as the framework for large-scale fault-tolerant quantum computers [19, 20, 21]. The ability to control and optimize such a multi-qubit array necessarily marks the next milestone in spin qubits.
@@
Analogous to how the basic instruction sets for different classical CPU architectures are different, quantum computers also have platform-dependent instruction sets. The general quantum computing theory favors controlled-not (CNOT) as the universal two-qubit gate [22]. Any algorithm may nevertheless be carried out with a different set of universal gates. For the best performance, a quantum algorithm should be compiled with the most natural gates for the physical platform it is applied upon. For spin qubits in particular, we consider a quantum gate as intrinsic if it can be attained with a single-step evolution under exchange interaction only. The controlled-phase (CPhase) or in particular controlled-Z (CZ) gate is an intrinsic two-qubit gate as it can be performed by simply turning on the exchange coupling [23, 24, 25]. In comparison, the CNOT gate for spin qubits is not intrinsic by our definition as it relies on external microwave drive to resonantly select the transition states [26]. In general, the “DC control” for CZ gate is less susceptible to noise compared with the “AC control” for CNOT [27, 28, 29], and is arguably easier to control and calibrate. Moreover, intrinsic gates need not be limited to two qubits. Most quantum algorithms require imposing joint transformations on multiple qubits. In these scenarios, intrinsic multi-qubit gates can often be used in place of equivalent clusters of single-qubit and two-qubit gates. Such replacements can boost efficiency, reducing control expenditures, while also less prone to errors [30]. Giving these obvious benefits, an abundance of recent studies have been devoted to multi-qubit gates in various systems such as Rydberg atoms [31, 32, 33, 34], superconducting qubits [30, 35, 36, 37] and trapped irons [38, 39]. For spin qubits, three-qubit Toffoli gate have been proposed and implemented using resonant microwave pulse [40, 41]. Notably, studies on linear quantum dot chains have revealed coherent “superexchange” oscillations of the boundary states [42, 43], and exhibit long-range coupling [44, 45]. These studies clearly demonstrate the possibility of multi-qubit gates with DC control and hint a larger class of intrinsic multi-qubit gates.
@@
Spin arrays coupled by exchange interaction have long fascinated theoretical investigations [46, 47, 48, 49, 50]. However, to-date it is still unclear what and how multi-qubit gates can be achieved in general. In this paper, we fill this important gap in current research with a comprehensive study on the qubit dynamics of quantum dot arrays. Extending the theoretical formalisms developed in an earlier paper [51], we greatly generalize the CZ/CPhase gate to a large class of multi-qubit gates intrinsically accessible to spin qubits. Furthermore, we also show these multi-qubit gates are scalable by proposing an efficient theoretical scheme to dynamically calibrate and control the inter-qubit couplings. While minimizing the crosstalk problem for selective controls, our scheme can also simultaneously protect the qubits from environmental couplings. Hopefully, our findings can pave the way for future large-scale quantum computing using spin qubits.
@@ II Dynamics of the dot array
Let us consider a spin-qubit chip defined in a semiconductor nanostructure using an array of quantum dots. An example set up is illustrated in Fig. 1(a) for a two-dimensional grid of quantum dots. The dots can also form a one-dimensional chain or even a three-dimensional lattice. Here we will use the general expression of “array” irrespective of the dimensionality. The dot array has suitable confining and chemical potentials that capture free charge carriers (electrons or holes) in the half-filling regime, i.e., only a single charge carrier occupies the lowest-energy orbit of each dot. A static magnetic field is applied to the system, lifting the two-fold spin degeneracy of the ground-level orbits and allowing localized spin qubits to be associated with the dots.
@@ II.1 Entangled state representation of the effective Hamiltonian
The Fermi-Hubbard model is a good starting point describing such quantum dot array [58, 59]. The low-energy states localized to all the dots can be normalized to form a basis $\{|\phi_{j,\sigma}\rangle\}$, where $j$ represents the dot location and $\sigma\in\{\uparrow,\downarrow\}$ is the spin index. Using the annihilation $a_{j\sigma}$ and creation $a^{\dagger}_{j\sigma}$ operators for the basis states, one can write down a second-quantized Hamiltonian for the dot array. We split this Hamiltonian into a dot part and a tunnel-coupling part $H=H_{\mathrm{d}}+H_{\mathrm{t}}$. The dot part describes the energy costs for filling 0 to 2 spins into the ground-level orbit to each dot,
$H_{\mathrm{d}}=\sum_{j\sigma}\left[\bigl{(}\mu_{j}+{\operatorname{sig}(\sigma_{j})}\,\tfrac{1}{2}\varepsilon_{\mathrm{Z},j}\bigr{)}n_{j\sigma}+\tfrac{1}{2}Un_{j\sigma}n_{j\bar{\sigma}}\right],$ (1)
@@
where $n_{j\sigma}=a^{\dagger}_{j\sigma}a_{j\sigma}$ is the number operator, $\mu_{j}$ and $\varepsilon_{\mathrm{Z},j}$ are the local chemical potential and the electron Zeeman energy at site $j$, the spin sign is defined by $\operatorname{sig}(\uparrow\downarrow)=\pm 1$ and $U$ is the energy cost for filling two antiparallel spins into the same orbit. The tunneling part is given by,
$H_{\mathrm{t}}=\sum_{\langle j,k\rangle}\sum_{\sigma}\left(t^{jk}_{\sigma\sigma}a^{\dagger}_{j\sigma}a_{k,\sigma}+s^{jk}_{\sigma\bar{\sigma}}a^{\dagger}_{j\sigma}a_{k,\bar{\sigma}}\right),$ (2)
where $t^{jk}_{\sigma\sigma}$ and $s^{jk}_{\sigma\bar{\sigma}}$ represent the spin dependent tunneling coefficients between adjacent dots $j,k$. We note that SOI mixes the orbit and the spin wave functions, while also opening a spin-flipping channel when tunneling among the dots [60].
Examining the spin-dependent tunneling process, for example, among dot $j$ and $k$. There are four spin conserved coefficients in addition to four spin flipped terms due to SOI and differences in spin axises. A key assumption for our theory is that the time reversal symmetry is approximately preserved for our system. This requires the Zeeman splitting energy to be much smaller than the characteristic barrier height between the dots. Hermicity of the Hamiltonian and well as time reversal symmetry implies that only two independent coefficients exists: the spin-conserved tunneling coefficient and spin-flipped tunneling coefficient,
$t_{j\downarrow,k\downarrow}=t_{k\uparrow,j\uparrow}=(t_{j\uparrow,k\uparrow})^{*}=(t_{k\downarrow,j\downarrow})^{*}$ $\propto t_{\langle{jk}\rangle},$ (3)
$s_{j\downarrow,k\uparrow}=s_{j\uparrow,k\downarrow}=(s_{k\downarrow,j\uparrow})^{*}=(s_{k\uparrow,j\downarrow})^{*}$ $\propto s_{\langle{jk}\rangle},$
@@
where we have introduced dimensionless coefficients for spin-conserved ($t$) and spin-flipped ($s$) tunneling. They satisfy the normalization condition $\lvert t\rvert^{2}+\lvert s\rvert^{2}=1$. In particular, in systems where direct-exchange dominates and $g$ factors are scalar, we can have the explicit parametrization $t=\cos(\gamma_{\mathrm{SO}})-\mkern 1.0mu\mathrm{i}\mkern 1.0mu\sin(\gamma_{\mathrm{SO}})\cos(\vartheta_{\mathrm{B}})$ and $s=-\mkern 1.0mu\mathrm{i}\mkern 1.0mu\sin(\gamma_{\mathrm{SO}})\sin(\vartheta_{\mathrm{B}})$ [51], where $\gamma_{\mathrm{SO}}$ and $\vartheta_{\mathrm{B}}$ characterize the SOI strength and the effective magnetic field angle. These relations make it possible to experimentally control the values of $t$ and $s$ by varying the SOI strength.
@@
The normalized tunneling coefficients uniquely determines an entangled state $|\xi_{w=\langle{j,k}\rangle}\rangle$ that lives on the four-dimensional Hilbert space defined by the half-filling states of the connecting vertices,
@@
$|\xi_{w}\rangle=\frac{1}{\sqrt{2}}\Bigl{(}s_{w}^{*}|\uparrow\uparrow\rangle_{w}+t_{w}^{*}|\uparrow\downarrow\rangle_{w}-t_{w}|\downarrow\uparrow\rangle_{w}+s_{w}|\downarrow\downarrow\rangle_{w}\Bigr{)}.$ (4)
@@
$H=H_{\mathrm{dot}}+H_{\mathrm{ex}}=\sum_{j}\frac{1}{2}\varepsilon^{\mathrm{Z}}_{j}\sigma_{j}^{\mathrm{Z}}-\sum_{\mathclap{w}}J_{w}|\xi_{w}\rangle\langle\xi_{w}|,$ (5)
where $H_{\mathrm{dot}}$ is the Hamiltonian of the quantum dots, which is the summation of the effective Zeeman splitting energies $\varepsilon^{\mathrm{Z}}_{j}$ along the Pauli operator $\sigma_{j}^{\mathrm{Z}}=|\uparrow\rangle\!\langle\uparrow|_{j}-|\downarrow\rangle\!\langle\downarrow|_{j}$ of all individual dots. The exchange term $H_{\mathrm{ex}}$ describes the coupling among connecting dots (edges). The coupling strength is specified by its exchange energy,
$J_{\langle j,k\rangle}=\frac{T_{jk}}{2}\left(\frac{1}{U-\mu_{j}+\mu_{k}}+\frac{1}{U-\mu_{k}+\mu_{j}}\right),$ (6)
which is a function of the tunneling amplitude, dot charging energy and the chemical potentials of the connecting dots. Equation (5) represents the effect of direct exchange in terms the entangled-state representation. We explicitly derive it in the Supplementary Material. Besides mathematically compact, the entangled-state representation of has many nice properties that we will take advantage of in the development of multi-qubit gates.

### v2

@@ Abstract
We study multi-qubit quantum gates intrinsic to an array of semiconductor quantum dots and investigate how they can be implemented in a scalable way. The intrinsic quantum gates refer to the class of natural-forming transformations in the qubit rotating-frame under direct exchange coupling, and can be recognized as an instruction set of spin-qubit chips. Adopting perturbative treatment, we can model the intrinsic gates by first-order dynamics in the coupling strength and develop a general formalism for identifying the multi-qubit intrinsic gates under arbitrary array connectivity. The advantageous applications of the intrinsic gates in quantum computing and quantum error correction are explored. Factors influencing the fidelities of the multi-qubit intrinsic gates are also discussed. To overcome the problem of inhomogeneous coupling, we propose a theoretical scheme in which single-qubit pulses are applied to dynamically calibrate the connecting bonds. This scheme can be further combined with periodic dynamical decoupling for robust implementations of multi-qubit gates in large-scale quantum computers.
@@ I Introduction
Semiconductor quantum dots are promising physical platforms for universal quantum computing [1, 2, 3, 4]. The spins of electrons (or holes) are natural two-level systems that can be selectively manipulated and brought into interactions by confining them with artificial structures. Spin qubits defined in quantum dots enjoy several unique advantages, including their miniature size and compatibility with modern semiconductor fabrication techniques, that make them suitable for creating large-scale quantum chips [5, 6]. Under intensive research efforts from the global community, the technologies and theories behind spin qubits have seen significant advancements over recent years. Continuous developments in key performance metrics such as the coherence times, operation frequencies and gate fidelities have been made [7]. Spin qubits in the state-of-the-art can be prepared and measured with fidelity exceeding $99\%$ [8], manipulated at frequency over 540 MHz [9] and work in temperature above 4K [10]. The fidelity figures for single-qubit and two-qubit gates have been pushed beyond the fault-tolerance threshold value [11, 12]. Industrial manufacturing has also been demonstrated for quantum dot arrays [13, 14].
@@
The natural roadmap of any quantum computing scheme is to incorporate gradually increasing quantities of qubits for realizing useful quantum algorithms. However, accurate control of multiple qubits can be a highly nontrivial task that implies much more than just putting the qubits together. For selective control of the interqubit coupling, the crosstalk effects of the control signals must be suppressed in an efficient manner. In addition, for robust implementation of multi-qubit gates, a method to overcome the inhomogeneous interqubit coupling is also indispensable. Here the key challenge is that the relevant control resources should scale well with the number of qubits, for potential application to large-scale chips with millions of qubits. For spin qubits, a viable way to scale up the qubit count is to employ an extensible two-dimensional array of quantum dots, as recently demonstrated for crossbar arrays of hole qubits [15, 16, 17, 18]. This configuration is also compatible with the surface code, a topological error correction code commonly conceived as the framework for large-scale fault-tolerant quantum computers [19, 20, 21]. The ability to control and optimize such a multi-qubit array necessarily marks the next milestone in spin qubits.
@@
The canonical quantum computing theory favors the controlled-not (CNOT) gate as the universal two-qubit gate [22]. Any algorithm may nevertheless be carried out with a different set of universal gates. Analogous to the distinctive basic instruction sets for different classical CPU architectures, quantum computers also have platform-dependent instruction sets. For optimal performance, a quantum algorithm should be compiled into the most natural gates for the physical platform it is applied upon. For spin qubits in particular, we consider a quantum gate as intrinsic if it can be attained with a single-step evolution for qubits coupled by exchange interaction only. The controlled-phase (CPhase) or in particular controlled-Z (CZ) gate is an intrinsic two-qubit gate as it can be performed by simply turning on the exchange coupling [23, 24, 25]. In comparison, the CNOT gate for spin qubits is not intrinsic by our definition as it relies on external microwave drive to resonantly select the transition states [26]. In general, the “DC control” for CZ gate is less susceptible to noise compared with the “AC control” for CNOT [27, 28, 29], and is arguably easier to control and calibrate. Moreover, intrinsic gates need not be limited to two qubits. Most quantum algorithms require imposing joint transformations on multiple qubits. In these scenarios, intrinsic multi-qubit gates can often be used in place of equivalent clusters of single-qubit and two-qubit gates. Such replacements can boost efficiency, reducing control expenditures, while also less prone to errors [30]. Giving these obvious benefits, an abundance of recent studies have been devoted to multi-qubit gates in various systems such as Rydberg atoms [31, 32, 33, 34], superconducting qubits [30, 35, 36, 37] and trapped irons [38, 39]. For spin qubits, the three-qubit Toffoli gate has been proposed and implemented using resonant microwave pulse [40, 41]. Notably, studies on linear quantum dot chains have revealed coherent “superexchange” oscillations of the boundary states [42, 43], and exhibit long-range coupling [44, 45]. These studies clearly demonstrate the possibility of multi-qubit gates with DC control and hint a larger class of intrinsic multi-qubit gates.
@@
Spin arrays coupled by exchange interaction have long fascinated theoretical investigations [46, 47, 48, 49, 50]. However, to-date it is still unclear what and how multi-qubit gates can be achieved in general. In this paper, we fill this important gap in current research with a comprehensive study on the qubit dynamics of quantum dot arrays. Extending the theoretical formalisms developed in an earlier paper [51], we greatly generalize the CZ/CPhase gate to a broad class of multi-qubit gates intrinsically accessible to spin qubits. Furthermore, we also show these multi-qubit gates are scalable by proposing an efficient theoretical scheme to dynamically calibrate and control the inter-qubit couplings. While minimizing the crosstalk problem for selective controls, our scheme can also simultaneously protect the qubits from environmental couplings. Hopefully, our findings can pave the way for future large-scale quantum computing using spin qubits.
@@ II Dynamics of the dot array
II.1 Entangled state representation of the effective Hamiltonian
The quantum chip for our study is a semiconducting nanostructure featuring an array of quantum dots. The dots can be typically arranged in 1 or 2 spatial dimensions and we use the word “array” irrespective of dimensionality. The dot array is kept at low temperature and has a suitable confining potential such that free charge carriers (electrons or holes) are captured in the half-filling regime, i.e., only a single carrier occupies each dot. To define spin qubits associated with the dots, a static magnetic field is applied to the system, lifting the spin degeneracy of the ground-level orbits. These dots can have different local Landé $g$-factor tensors, hence the spatial quantization axes and Zeeman splitting energies for each dot can differ. This device also possesses considerable spin-orbital interaction (SOI) to enable fast manipulation of spin states with electric signals [55]. The SOI can be either intrinsic to the nanostructure [56] or artificially introduced by slanting magnetic field [57]. An example set up is illustrated in Fig. 1(a) for a two-dimensional grid of quantum dots.
@@ II.1 Entangled state representation of the effective Hamiltonian
The Fermi-Hubbard model is a good starting point for describing such quantum dot array [58, 59]. The low-energy states localized to all the dots can be normalized to form a basis $\{|\Phi_{j\sigma}\rangle\}$, where $j$ represents the dot location and $\sigma\in\{\uparrow,\downarrow\}$ is the spin index. Introducing the annihilation $a_{j\sigma}$ and creation $a^{\dagger}_{j\sigma}$ operators for the basis states, one can write down a second-quantized Hamiltonian that consists of a “dot” part $H_{\mathrm{dot}}$ and a “tunneling” part $H_{\mathrm{tun}}$. The dot part describes the energy cost for filling 0 to 2 charge carriers into the ground-level orbit of each quantum dot,
$H_{\mathrm{dot}}=\sum_{j\sigma}\Bigl{[}\bigl{(}\mu_{j}+{\operatorname{sig}(\sigma_{j})}\,\frac{1}{2}\varepsilon^{\mathrm{Z}}_{j}\bigr{)}n_{j\sigma}+\frac{1}{2}U_{j}n_{j\sigma}n_{j\bar{\sigma}}\Bigr{]},$ (1)
@@
where $n_{j\sigma}=a^{\dagger}_{j\sigma}a_{j\sigma}$ is the number operator for the $\sigma$ spin state at dot $j$, $\mu_{j}$ and $\varepsilon^{\mathrm{Z}}_{j}$ are the local chemical potential and the Zeeman splitting energy at dot $j$, the spin sign is defined by $\operatorname{sig}(\uparrow\downarrow)=\pm 1$ and $U_{j}$ is the Coulomb charging energy for filling two antiparallel spins into the same dot $j$. Interdot tunneling of charge carriers is described by
$H_{\mathrm{tun}}=\sum_{\langle j,k\rangle}\sum_{\sigma}\left(t^{jk}_{\sigma\sigma}a^{\dagger}_{j\sigma}a_{k,\sigma}+t^{jk}_{\sigma\bar{\sigma}}a^{\dagger}_{j\sigma}a_{k,\bar{\sigma}}\right),$ (2)
where the first summation is over all pairs of neighboring dots $\langle j,k\rangle$ that can be coupled via electronic tunneling, $t^{jk}_{\sigma\sigma}$ and $t^{jk}_{\sigma\bar{\sigma}}$ represent the spin dependent tunneling coefficients between adjacent dots $j,k$. Examining the tunneling process among a pair of dots $j$ and $k$, we can identify four spin-conserved terms in addition to four spin-flipped terms. The spin-flipping can be attributed to differences in the local quantization axes, and/or the SOI effects of the structure [60, 61].
A necessary assumption for our theory is that the typical Zeeman energy associated with each dot is much smaller than the potential barrier between the dots. As explained in the Appendix A.2, this assumption translates to a weak time reversal symmetry condition that allows the tunneling coefficients to be reduced into only two complex numbers,
$t^{jk}_{\uparrow\uparrow}$ $=(t^{jk}_{\downarrow\downarrow})^{*}=t^{kj}_{\downarrow\downarrow}=(t^{kj}_{\uparrow\uparrow})^{*}$ $\equiv\,t_{jk}\;\propto\,\tilde{t}_{jk}$ (3)
$t^{jk}_{\uparrow\downarrow}$ $=-(t^{jk}_{\downarrow\uparrow})^{*}=-t^{kj}_{\uparrow\downarrow}=(t^{kj}_{\downarrow\uparrow})^{*}$ $\equiv\,s_{jk}\;\propto\,\tilde{s}_{jk}.$
@@
where we have introduced the spin-conserved ($t_{jk}$) and the spin-flipped ($s_{jk}$) tunneling coefficients. This relation among tunneling coefficients is illustrated in Fig. 1(b). For convenience, we also introduce the dimensionless tunneling coefficients $\tilde{t}_{jk}$ and $\tilde{s}_{jk}$, that are normalized to unity $\lvert\tilde{t}_{jk}\rvert^{2}+\lvert\tilde{s}_{jk}\rvert^{2}=1$. In particular, for systems where the $g$-factors are scalar, we can further derive the explicit parametrization $\tilde{t}=\cos(\gamma_{\mathrm{SO}})+\mkern 1.0mu\mathrm{i}\mkern 1.0mu\sin(\gamma_{\mathrm{SO}})\cos(\vartheta_{\mathrm{B}})$ and $\tilde{s}=\mkern 1.0mu\mathrm{i}\mkern 1.0mu\sin(\gamma_{\mathrm{SO}})\sin(\vartheta_{\mathrm{B}})$, where $\gamma_{\mathrm{SO}}$ and $\vartheta_{\mathrm{B}}$ characterize the SOI strength and the effective magnetic field angle in the vicinity of a pair of dots (dot-subscripts omitted here) [51]. These relations suggest it is possible to experimentally control the portion of spin-conserved and spin-flipped coefficients by varying the SOI strength and/or magnetic field angle.
The immediate state after tunneling involves two spins occupying the same dot and is not energetically favored due to the large Coulomb repulsion energy. The theory of quantum mechanics, however, allows brief transition to such high energy states. This virtual tunneling process gives rise to direct exchange interaction, assumed to be the main interqubit coupling mechanism for our device. To specify the dynamics of the half-filling states of interest, we introduce the computational basis states for an array of $N_{Q}$ quantum dots,
$|n\rangle\equiv|\sigma_{n_{1}},\sigma_{n_{2}},\cdots,\sigma_{n_{N_{Q}}}\rangle,\quad\text{for }n=1,2,\cdots 2^{N_{Q}},$ (4)
where the component $\sigma_{n_{j}}\in\{\uparrow,\downarrow\}$ is the spin state at dot $j$ and $n_{j}$ can be viewed as the binary digit of $n$. Each $|n\rangle$ is an Fermionic (antisymmetric) multi-body wave function and Eq. (4) can be viewed as a Fock space representation from a formal point of view.
@@
The tunneling coefficients for a pair of connected dots $w=\langle{j,k}\rangle$ can be used to specify an entangled state that lives on the four-dimensional space spanned by the half-filling states of the connecting dots,
@@
$|\xi_{w}\rangle=\frac{1}{\sqrt{2}}\Bigl{(}\tilde{s}_{w}|\uparrow\uparrow\rangle_{w}-\tilde{t}_{w}|\uparrow\downarrow\rangle_{w}+\tilde{t}^{*}_{w}|\downarrow\uparrow\rangle_{w}+\tilde{s}^{*}_{w}|\downarrow\downarrow\rangle_{w}\Bigr{)}.$ (5)
@@
$H=\underbrace{\sum_{j}\frac{1}{2}\varepsilon^{\mathrm{Z}}_{j}\sigma_{j}^{\mathrm{Z}}}_{H_{\mathrm{qu}}}\quad-\underbrace{\sum_{\mathclap{w=\langle{j,k}\rangle}}J_{w}|\xi_{w}\rangle\langle\xi_{w}|}_{H_{\mathrm{ex}}},$ (6)
where the first part $H_{\mathrm{qu}}$ defines the energy splitting of the qubits, which is the summation of the effective Zeeman splitting energies $\varepsilon^{\mathrm{Z}}_{j}$ along the Pauli operator $\sigma_{j}^{\mathrm{Z}}=|\uparrow\rangle\!\langle\uparrow|_{j}-|\downarrow\rangle\!\langle\downarrow|_{j}$ for all dots. The exchange term $H_{\mathrm{ex}}$ describes the coupling among connecting dots (edges/bonds), with coupling strength specified by the exchange energy,
$J_{jk}=\frac{\lvert t_{jk}\rvert^{2}+\lvert s_{jk}^{2}\rvert}{2}\left(\frac{1}{U_{j}+\mu_{j}-\mu_{k}}+\frac{1}{U_{k}+\mu_{k}-\mu_{j}}\right).$ (7)
We provide an explicitly derivation of this result in Appendix A. However, one can show that these two representations are equivalent.
Equation (6) represents the effect of direct exchange in terms entangled states that live on connected edges. This is quite distinctive from the traditional representation of the exchange interaction $H_{\mathrm{ex}}=\bm{S}_{j}\mathcal{J}_{jk}\bm{S}_{k}$, where $\mathcal{J}_{jk}$ is generally a tensor. Besides being mathematically compact, our entangled-state representation allows easier treatment of the anisotropic exchange and has many nice properties that we will take advantage of in the development of multi-qubit gates.